OpenStudy (chaise):

Can anyone help me implicitly differentiate this?? X^3+xtan^-1(y)=e^y

OpenStudy (anonymous):

What is giving you a problem. Use normal differentiation rules and remember chain rule: [g(y)]'=g'(y)y'

OpenStudy (chaise):

\[x^3+x \tan^{-1} (y)=e^y\] Differentiating I got: 3x^2+x/(1+y^2)(dy/dx)+tan^(-1)(y)=e^y(dy/dx)

OpenStudy (chaise):

Then to simplify: 3x^2+tan^(-1)(y)=e^y(dy/dx)-x/(1+y^2)(dy/dx) Factoring out dy/dx of RHS 3x^2+tan^(-1)(y)=dy/dx(e^y-(x/(1+y^2)) Solving for dy/dx 3x^2+tan^(-1)(y)/(e^y-x/(1+y^2)=dy/dx

OpenStudy (anonymous):

sorry meant the first one from left to right

OpenStudy (chaise):

That is what I am having trouble with - why should it be in the numerator? dy/dx of arctan(x) = 1/(x^2+1) Right?

OpenStudy (anonymous):

Let g(y)=arctan(y) We know [g(y)]'=g(y)*(dy/dx) by chain rule. So we have \[\frac{dy}{dx}\frac{1}{1+y^{2}}\] The dy/dx is technically in the numerator of the fraction

OpenStudy (anonymous):

Sorry [g(y)]'=g'(y)*(dy/dx)

OpenStudy (chaise):

So this is what it should look like? (dy/dx)/1/(1+y^2) (dy/dx)*(1+y^2)/1 (1+y^2)dy/dx (not forgetting to multiply by x in my final answer because of the product rule)

OpenStudy (anonymous):

No no you don't need to manipulate it. The derivative of arctan(y) is just as show above